# TriSAM: Tri-Plane SAM for zero-shot cortical blood vessel segmentation in VEM images

### 25 Jan 2024 v1

@@ 3.2 TriSAM 3D Segmentation
Recursive Redirection. SAM-based tracking is simple and fast for short-term tracking. But it may fail when shape and scale change dramatically as shown in Figure 3c. To enable long-term tracking, we propose a recursive redirection method to exploit the 3D blood vessel structure, which keeps track of potential turning points in blood vessels. Specifically, for each seed, we maintain a turning point tree where the root is the initial seed. In each iteration, we sample a left node in the turning point tree and perform Tri-Plane selection and track along the chosen plane. As blood vessels may change directions within this iteration, we then split this node to obtain new potential turning points. We find that the non-best planes potentially identify turning points or endpoints accurately as shown in Figure 4, Therefore, we propose to generate the turning points according to the segmentation results at non-best planes. In particular, when tracking along the z-axis, we apply SAM to segment blood vessels in the zx and zy planes. Potential turning points are identified as points in these segments with the smallest and largest z-values.
4 Experiments
@@
4.1 Experimental Settings and Implementation Details
Dataset Split. We used the whole dataset for evaluation. For the supervised setting, we randomly sampled subvolumes are annotated. We sampled 4 subvolumes of size 256$\times$256$\times$256 for the BvEM-Macaque dataset amounting to around 10% of the dataset, and 12 subvolumes of size 384$\times$384$\times$384 for each of the BvEM-Mouse and BvEM-Human datasets amounting to around 1% of each dataset. These annotated subvolumes were divided into a 3-1 train-valid split.
Implementation Details. Since the volume is too large, non-overlapping subvolumes are used for segmentation and the results are fused to form the final prediction. For BvEM-Macaque, we directly predict the whole volume. For others, $k\times 1024\times 1024$ subvolumes are used since $1024\times 1024$ is the default resolution for SAM where $k=661$ for BvEM-Human and $k=818$ for BvEM-Mouse. For preprocessing, we utilize defliker method and temporal smoothing on the z-axis to remove the artifacts and missing slices. For postprocessing, holes are filled and small connected components are removed after the segmentation. Global color thresholding is used to generate the initial seeds and $\eta$ is set to 0.98 as shown Figure 7 (left). Given the significantly greater complexity of human blood vessels, we employ the manually generated seeds for the challenging BvEM-Human dataset. For SAM-based tracking, the probability threshold $\tau$ is set to 0.8 according to the experiment shown in Figure 7 (right). Unless explicitly stated otherwise, we use MobileSAM (Zhang et al., 2023) instead of the standard SAM (Kirillov et al., 2023) in all conducted experiments to improve the inference speed. All experiments are conducted on an NVIDIA-A100 GPU.
Evaluation Metrics. Following Weigert et al. (2020), we use Precision, Recall, and Accuracy as the metrics to evaluate the performance:
$Precision=\frac{TP}{TP+FP},\quad Recall=\frac{TP}{TP+FN},\quad Accuracy=\frac{TP}{TP+FP+FN},$ (1)
where $TP$, $FP$, and $FN$ are instance-level true positive, false positive, and false negative respectively. We use instance-level metrics since it is more sensitive to split errors. In particular, the Hungarian algorithm is used to match ground-truth instances and predicted instances with negative Accuracy as the cost matrix.
Table 2: Benchmark results on the proposed BvEM dataset. We evaluate the initial blood vessel annotation to show the amount of proofreading effort. We compare the proposed TriSAM with various baseline methods under fully supervised with limited training data, unsupervised, and zero-shot settings.
Method Training Data BvEM-Mouse BvEM-Macaque BvEM-Human
@@ 4.1 Experimental Settings and Implementation Details
Initial Annotation  [^1] Results from original dataset papers. BvEM-Mouse and BvEM-Macaque are automatically segmented while BvEM-Human is manually labeled. Unknown 93.74 36.62 35.74 1.65 23.17 1.57 100.00 25.68 25.68
3D UNet Limited 13.45 0.91 0.67 16.04 89.65 15.75 68.63 2.46 2.43
nnUNET Limited 3.54 49.55 6.59 24.34 29.57 23.74 3.44 23.20 5.86
Color Thresholding None 86.45 37.32 35.26 95.14 21.65 21.42 41.77 1.92 1.87
MAESTER None 2.16 18.95 .94 22.08 40.30 16.64 0.29 5.03 0.27
SAM + IoU Tracking External 63.59 0.27 0.27 74.39 1.89 1.88 18.19 23.58 12.92
TriSAM External 84.12 66.75 59.28 78.41 74.97 62.14 31.35 25.57 16.39
@@
Figure 5: Instance segmentation results of comparison methods on BvEM-Macaque. Different colors indicate different instances. Color thresholding and 3D UNet often produce false positives, whereas SAM + IoU tracking tends to miss a significant portion of blood vessels. Among the comparison methods, TriSAM segmentation stands out as the most effective.
4.2 Benchmark Results
@@
Methods in Comparison. We compare the proposed TriSAM with both zero-shot baselines and supervised methods. The compared zero-shot methods are global color thresholding and SAM + IoU tracking. For color thresholding, we first perform (3D) Gaussian blurring with $\sigma=1$ on 3D chunks (10x512x512). Then we label all voxels that are 3 standard derivation above mean as positive. Finally, connected components with less than 1000 voxels are filtered out. For SAM + IoU Tracking, we segment all objects in each z-slice of the dataset using automatic mask generation. Then we track each blood vessel using the first labelled slice as seeds. Our simple tracking algorithm finds the mask in next slice with maximum IoU with the current slice. If the max IoU is above a threshold, we assign this mask to the current object and continue tracking. We also tried SAM + IoU tracking with microSAM Archit et al. (2023b) weights that have been finetuned on EM images. This model however does not work well on our datasets. We expect this is because microSAM has been finetuned on high-resolution EM images and does not generalize to our low-resolution datasets. We further compare TriSAM with the supervised method 3D U-Net (The MICrONS Consortium et al., 2023). We use the implementation from Wolny et al. (2020). Individual models are trained on $5\%$ of each dataset. We use $64\times 192\times 192$ dimensional cubes for training and train for $20000$ iterations.
@@
Results Analysis. The experimental results are shown in Table 2. First, both the Color Thresholding and SAM + IoU Tracking methods exhibit significant performance variability across various sub-datasets, highlighting the diversity of our dataset and the sensitivity of these methods to different species. Furthermore, both of these unsupervised methods demonstrate relatively poor performance, underscoring the challenges of the zero-shot setting in the BvEM dataset. Additionally, the 3D UNet, as a supervised learning approach, also yields very subpar results, indicating poor generalization of models trained with limited data. Finally, TriSAM outperforms the comparison methods by a large margin as it not only accurately segments the boundary but also tracks the blood vessels in the long term. Note that the results from the original dataset paper are also reported for reference.
@@
Qualitative Results. The final instance segmentation results on BvEM-Macaque are shown in Figure 5. Color thresholding segments bright pixels, inadvertently capturing nuclei cells while overlooking darker pixels corresponding to blood vessels. Training the 3D UNet model with limited data results in confusion with background elements. IoU tracking fails to capture a significant portion of the blood vessel, revealing its ineffectiveness in tracking. TriSAM prediction emerges as the most accurate method, affirming the effectiveness of the proposed approach.
4.3 Ablation Study
@@
Table 3: Ablation study results on different plan selection strategies. The proposed Tri-plane approach achieves the best overall accuracy with comparable speed.
Method Precision (%) Recall (%) Accuracy (%) Speed (sec)
@@
Figure 6: Comparison of segmentation results from one seed using different plane selection strategies.
@@
Figure 6: Comparison of segmentation results from one seed using different plane selection strategies.
Figure 7: Ablation study results on the hyperparameters $\eta$ and $\tau$.
Effect of the Tri-Plane Selection. We first compare our method with the single-plane methods to evaluate the effectiveness of Tri-Plane. For the single-plane method, we choose one plane as the main plane and only track along the chosen plane. The results on the BvEM-Macaque dataset are shown in Tabel 3. We have observed significant differences in performance among the three single-plane methods, with accuracy ranging from 11.53% to 50.75%, depending on the chosen tracking plane, which indicates the importance of the chosen tracking plane. This variability can be attributed to the tubular nature of blood vessel extensions within biological organisms, resulting in the generation of intricate mask shapes in certain planes, while simpler mask shapes are produced in others. Then we fuse the results of three Single-Plane methods and attain a higher accuracy of 57.41%, which demonstrates that the segmentation results from different planes exhibit a high degree of complementarity with each other. Instead, Tri-Plane exploits the blood vessel 3D structures by tracking the blood vessels along a suitable plane and attains the highest accuracy of 62.14%.
We visualize the segmentation results with one initial seed on BvEM-Macaque in Figure 7. We see that the performance is sensitive to the selection of the tracking plane. If the plane is not well-selected, the segmentation result can be empty as the example of tracking along the x-axis shows. The best result is tracking along the y-axis. However, it still falls short of the proposed method’s performance, as the latter takes into account potential turning points to exploit the 3D structure and perform long-term tracking.
Table 4: Ablation study results on different redirection strategies. The proposed recursive redirection approach achieves the best overall accuracy with comparable speed.
Strategy Precision (%) Recall (%) Accuracy (%) Speed (sec)
@@
Dense Redirection 86.94 24.84 23.95 2001($\uparrow$ 552%)
Recursive Redirection 78.41 74.97 62.14 335 ($\uparrow$ 9%)
@@
Effect of the Recursive Redirection Strategy. To perform long-term tracking and fully leverage the 3D blood vessel structure, we introduced recursive redirection by considering potential turning points. To validate its effectiveness, we report the results on the BvEM-Macaque dataset in Table 4, where the runtime comparison for segmentation prediction on the entire BvEM-Macaque data is also included. One naive baseline is to remove the recursive redirection component and not consider any potential turning points. This strategy is simple and fast but it fails to exploit the 3D shape prior leading to poor performance. Another strategy is to select the best plane for every tracking step/slice, which densely performs SAM segmentation on each step across three planes. Unfortunately, it significantly increases the computation cost. However, we were surprised to observe that the performance of the Dense Redirection strategy was even worse. This could be attributed to the frequent axis changes potentially leading to the omission of certain parts of the blood vessel and causing splitting errors.
@@
Table 5: Ablation study results on the choice of SAM models. The MobileSAM model achieves better performance with faster inference speed and smaller model size.
Method Pre (%) Rec (%) Acc (%) Speed Parameters
Ours + SAM (Kirillov et al., 2023) 77.65 66.03 55.48 1535s 615M
Ours + MobileSAM (Zhang et al., 2023) 78.41 74.97 62.14 335s 9.66M
Effect of the Mobile-SAM Model. To further improve the efficiency, we use MobileSAM instead of the original SAM in practice. As shown in Table 5, the inference time of MobileSAM is 22% of the original SAM which confirms that MobileSAM significantly improves the inference speed. Moreover, the performance of MobileSAM is even better than the original SAM, possibly because the distilled small model is less prone to overfitting to the original natural image domain.
@@
Quality Analysis We delve deeper into the plane dynamics in Figures 10 and 10. The majority of the selected plane predominantly tracks along the y-axis. This observation aligns with the experimental results presented in Table3, where it is evident that a single plane tracking with the y-axis outperforms the z-axis and x-axis. This is because the blood vessel flows mainly along the y-axis as shown in Figure 7. In Figure 10, we explore the size dynamics using various methods. The size variation observed with our proposed method is relatively smaller compared to tracking along the y-axis and significantly less than when tracking along the z-axis and x-axis.
Figure 8: Histogram of tracking planes.
Figure 9: Dynamics of tracking planes.
Figure 8: Histogram of tracking planes.
Figure 9: Dynamics of tracking planes.
Figure 10: Dynamics of size during tracking. (X) indicates the mean derivation is X.
5 Related Work
Blood Vessel Segmentation. Most existing VEM image segmentation algorithms were developed for neurons (Lee et al., 2019) and synapses (Wu et al., 2023a; Turner et al., 2020; Buhmann et al., 2021). The traditional VEM dataset size is too small to study blood vessel architecture. Recently, with rapid technology improvement, VEM sample size reached the cubic millimeter scale covering all the layers of cerebral cortex (Shapson-Coe et al., 2021; The MICrONS Consortium et al., 2023). This marks a significant milestone in connectomics research. The blood vessels were segmented manually in the human cortex (Shapson-Coe et al., 2021) and automatically using fully convolutional neural network (FCN) Tetteh et al. (2020) or 3D U-Net (The MICrONS Consortium et al., 2023; Livne et al., 2019; Çiçek et al., 2016). Due to the burden of annotation, efforts have been made to decrease the need for annotations (Dang et al., 2022). For images acquired using other related imaging methods, such as light microscopy and MRI, a variety of methods were developed and covered by some reviews (Goni et al., 2022; Moccia et al., 2018). Briefly, the blood vessels could be enhanced based on the nature of tubeness using filters, such as diffusion filtering and ridge filtering. However, the segmentation of cortical blood vessels in VEM images is seldom explored, mainly because there is a lack of publicly available benchmarks.
Evaluation Metrics. Quantitative evaluation is normally conducted by comparing with manual annotations that are used as the ground truth. The evaluation metrics could be classified into two categories: pixel-based and skeleton-based (Moccia et al., 2018). Pixel-based metrics measure the difference between segmentation and ground truth at the pixel or voxel level, such as Accuracy and Dice Similarity Coefficient (Dice, 1945). Skeleton-based metrics measure the difference of skeletons or centerlines, such as Overlap and Overlap Until First Error (Metz et al., 2008). Skeleton-based metric requires the blood vessel skeleton that might not always be available. It also requires the computation of point-to-point correspondence between the ground truth and computed skeleltons. All existing evaluation metrics are developed using datasets acquired using other imaging methods and we did not find any specific metric for VEM datasets. Here we adopt the metrics Precision, Recall, and Accuracy as we are measuring instance-level pixel prediction.
Segment Anything-Based Models. As a foundation model for image segmentation, the recently proposed Segment Anything Model (Kirillov et al., 2023) has garnered significant attention. Due to its exceptional zero-shot performance, SAM has been extended to a variety of domains including object tracking (Cheng et al., 2023b; Yang et al., 2023), image inpainting (Yu et al., 2023), image mattting (Yao et al., 2023), super-resolution (Lu et al., 2023), 3D point cloud (Liu et al., 2023), and image editting (Gao et al., 2023). Despite SAM’s remarkable generalization capabilities, it still encounters several challenges in practical applications. One of these challenges is the huge computation costs due to the heavyweight image encoder. FastSAM (Zhao et al., 2023) adopted a conventional CNN detector with an instance segmentation branch for the segment anything task with real-time speed. MobileSAM (Zhang et al., 2023) proposed decoupled distillation to obtain a small image encoder, which achieved approximately five times faster speed compared to FastSAM while also being seven times smaller in size. Therefore the MobileSAM is employed in our proposed method. Another challenge lies in the unsatisfactory performance of SAM when confronted with special domains, such as medical (Chen et al., 2023) or biological (Archit et al., 2023a) images, particularly in the context of 3D data. Deng et al. (2023) assessed the SAM model’s zero-shot segmentation performance in the context of digital pathology and showed scenarios where SAM encounters difficulties. Mazurowski et al. (2023) extensively evaluates the SAM for medical image segmentation across 19 diverse datasets, highlighting SAM’s performance variability based on prompts and dataset characteristics. To address the domain gap between natural and medical images, SAM-Adapter (Chen et al., 2023), SAM-Med2D (Cheng et al., 2023a), and Medical SAM Adapter (Wu et al., 2023b) introduced Adapter modules and trained the Adapter with medical images. They attained good performance on various medical image segmentation tasks. MedSAM (Ma & Wang, 2023) adapted SAM with more than one million medical image-mask pairs and attained accurate segmentation results. MicroSAM (Archit et al., 2023a) also presented a segment anything model for microscopy by fine-tuning SAM with microscopy data. Differing from these approaches that require model fine-tuning for adaptation, our method lifts the blood vessel segmentation capabilities of SAM from 2D images to 3D volumes without any model fine-tuning.
6 Conclusion
In this paper, we have contributed the largest-to-date public benchmark, the BvEM dataset, for cortical blood vessel segmentation in 3D VEM images. By addressing a significant gap in the field of neuroimaging, we have laid the foundation for advancing the understanding of cerebral vasculature at the microscale and its intricate relationship with neural function. We also developed a zero-shot blood vessel segmentation method, TriSAM, based on the powerful SAM model, offering an efficient and accurate approach for segmenting blood vessels in VEM images. With Tri-Plane selection, SAM-based tracking, and recursive redirection, our TriSAM effectively exploits the 3D blood vessel structure and attains superior performance compared with existing zero-shot and supervised technologies on BvEM across three species, marking a critical step towards unlocking the mysteries of neurovascular coupling and its implications for brain health and pathology. With the availability of the BvEM dataset and the TriSAM method, researchers are now equipped with valuable tools to drive breakthroughs in VEM-based cortical blood vessel segmentation and further our understanding of the brain’s intricate vascular network.
Appendix A Appendix
A.1 Failure Cases

### 09 Apr 2024 v2

@@ IV-D Recursive Seed Sampling
In practice, we maintain a tree of seed points where the root is the initial seed. In each iteration, we sample a left node in the tree, mark it as visited, and run TriSAM. Empirically, we find the non-best planes potentially identify turning points or endpoints accurately. Therefore, we propose to generate the turning points according to the segmentation results at non-best planes. Potential turning points are identified as points in these segments with the smallest and largest z-values. The algorithm of TriSAM is shown in Algorithm 1.
Algorithm 1 TriSAM algorithm
[…]
Initialize the seed tree $SeedTree=s$
for ${\textnormal{t}}_{i}\quad in\quad\operatorname{RemainNode}(SeedTree)$ do
[…]
     if $\operatorname{Valid}(Seeds)$ then
         $SeedTree=\operatorname{LeafSplit}(SeedTree,{\textnormal{t}}_{i},Seeds)$triangleright Grow the seed tree
     end if
end for
[…]
V Experiments
@@
V-A Experimental Settings
Implementation Details. Since the volume is too large, non-overlapping subvolumes are used for segmentation and the results are fused to form the final prediction. Initial seeds can be effectively generated with global color thresholding since the pixels of blood vessels are brighter than the background. Therefore, pixels darker than a threshold are considered as background while others are seed candidates. To improve efficiency, we only keep the center of each connected component as the final seeds. For BvEM-Macaque, we directly predict the whole volume. For others, $k\times 1024\times 1024$ subvolumes are used since it is the default resolution for SAM where $k=661$ for BvEM-Human and $k=818$ for BvEM-Mouse. For preprocessing, we utilize deflickering method and temporal smoothing on the z-axis to remove the artifacts and missing slices. For postprocessing, holes are filled and small connected components are removed after the segmentation. Global color thresholding is used to generate the initial seeds and $\eta$ is set to 0.98 as shown Fig. 8 (left). Given the significantly greater complexity of human blood vessels, we employ the manually generated seeds for the challenging BvEM-Human volume. For SAM-based tracking, the probability threshold $\tau$ is set to 0.8. according to the experiment shown in Fig. 8 (right). Unless explicitly stated otherwise, we use MobileSAM [36] instead of the standard SAM [15] in all conducted experiments to improve the inference speed. All experiments are conducted on an NVIDIA-A100 GPU.
Evaluation Metrics. We use the Precision, Recall, and Accuracy metric defined in [46], where $Accuracy=\frac{TP}{TP+FP+FN}$.
$Precision$ $=\frac{TP}{TP+FP},$ (1)
$Recall$ $=\frac{TP}{TP+FN},$ (2)
$Accuracy$ $=\frac{TP}{TP+FP+FN},$ (3)
where $TP$, $FP$, and $FN$ are instance-level true positive, false positive, and false negative respectively. We use instance-level metrics since it is more sensitive to split errors. In particular, the Hungarian algorithm is used to match ground-truth instances and predicted instances with negative Accuracy as the cost matrix. We used the whole dataset for evaluation and computed the score on the largest instance segment of each volume due to its dominating size.
TABLE II: Benchmark results on the proposed BvEM dataset. We evaluate the initial blood vessel annotation to show the significant amount of our proofreading effort.
Method Setting BvEM-Mouse BvEM-Macaque BvEM-Human
@@ V-A Experimental Settings
Initial annotation N/A 93.74 36.62 35.74 1.65 23.17 1.57 100.00 25.68 25.68
Color thresholding N/A 86.45 37.32 35.26 95.14 21.65 21.42 41.77 1.92 1.87
MAESTER [47] unsupervised 2.16 18.95 0.94 22.08 40.30 16.64 0.29 5.03 0.27
3D UNet [11] supervised 13.45 0.91 0.67 16.56 86.95 16.16 68.63 2.46 2.43
nnUNET [48] supervised 3.54 49.55 6.59 24.34 29.57 23.74 3.44 23.20 5.86
SAM+IoU tracking zero-shot 63.59 0.27 0.27 74.39 1.89 1.88 18.19 23.58 12.92
TriSAM (ours) zero-shot 84.12 66.75 59.28 78.41 74.97 62.14 31.35 25.57 16.39
@@
Figure 6: Qualitative instance segmentation results on the BvEM-Macaque volume. Different colors indicate different instances. Color thresholding and 3D UNet often produce false positives, whereas SAM+IoU tracking tends to miss a significant portion of blood vessels. Among the comparison methods, TriSAM segmentation stands out as the most effective.
V-B Benchmark Results
@@
Methods in Comparison. We compare the proposed TriSAM with both zero-shot baselines and supervised methods. The compared zero-shot methods are global color thresholding and SAM+IoU tracking. For color thresholding, we first perform (3D) Gaussian blurring with $\sigma=1$ on 3D chunks (10x512x512). Then we label all voxels that are 3 standard derivations above mean as positive. Finally, connected components with less than 1000 voxels are filtered out. For SAM+IoU Tracking, we segment all objects in each z-slice of the dataset using automatic mask generation. Then we track each blood vessel using the first labeled slice as seeds. Our simple tracking algorithm finds the mask in the next slice with maximum IoU with the current slice. If the max IoU is above a threshold, we assign this mask to the current object and continue tracking. We also tried SAM+IoU tracking with microSAM [49] weights that have been finetuned on EM images. This model however does not work well on our dataset. We expect this is because microSAM has been finetuned on high-resolution EM images and does not generalize to our low-resolution dataset. We further compare TriSAM with the supervised method 3D U-Net [11]. We use the implementation from [50]. For supervised methods, we cropped subvolumes at the center of each volume composing approximately 10% in size. These annotated subvolumes were divided into a 1-1 train-val split.
@@
Results Analysis. The results are shown in Tab. II. First, both the Color Thresholding and SAM+IoU Tracking methods exhibit significant performance variability across three volumes, highlighting the diversity of our dataset and the sensitivity of these methods to different species. Furthermore, both of these unsupervised methods demonstrate relatively poor performance, underscoring the challenges of the zero-shot setting in the BvEM dataset. Additionally, the 3D UNet, as a supervised learning approach, also yields subpar results, indicating poor generalization of models trained with limited data. Finally, TriSAM significantly outperforms other methods as it not only accurately segments the boundary but also tracks the blood vessels in the long term.
@@
Qualitative Results. The final instance segmentation results on BvEM-Macaque are shown in Fig. 6. Color thresholding segments bright pixels, inadvertently capturing nuclei cells while overlooking darker pixels corresponding to blood vessels. Training the 3D UNet model with limited data results in confusion with background elements. IoU tracking fails to capture a significant portion of the blood vessel, revealing its ineffectiveness in tracking. TriSAM prediction emerges as the most accurate method, affirming its effectiveness.
V-C Ablation Studies
@@
Effect of the Tri-plane Selection. We first compare our method with the single-plane methods to evaluate the effectiveness of the tri-plane. For the single-plane method, we choose one plane as the main plane and only track along the chosen plane. The results on the BvEM-Macaque volume are shown in Tab. III. We have observed significant differences in performance among the three single-plane methods, with accuracy ranging from 11.53% to 50.75%, depending on the chosen tracking plane, which indicates the importance of the chosen tracking plane. This variability can be attributed to the tubular nature of blood vessel extensions within biological organisms, resulting in the generation of intricate mask shapes in certain planes, while simpler mask shapes are produced in others. Then we fuse the results of three Single-Plane methods and attain a higher accuracy of 57.41%, which demonstrates that the segmentation results from different planes exhibit a high degree of complementarity with each other. Instead, We see that tri-plane selection exploits the blood vessel 3D structures by tracking along a suitable plane and attains the highest accuracy of 62.14%.
TABLE III: Ablation studies on different plane selection strategies. The proposed tri-plane approach achieves the best overall accuracy with comparable speed.
Method Pre. (%) Rec. (%) Acc. (%) Speed (sec)
@@
Figure 7: Segmentation results with different plane selection strategies.
@@
Figure 7: Segmentation results with different plane selection strategies.
Figure 8: Ablation study results on hyperparameters $\eta$ and $\tau$.
Figure 9: Histogram of the tri-plane selection.
Figure 10: Dynamics of the tri-plane selection.
Figure 9: Histogram of the tri-plane selection.
Figure 10: Dynamics of the tri-plane selection.
Figure 11: Dynamics of segment size.
We visualize the segmentation results with one initial seed on BvEM-Macaque in Fig. 8. The origin point is the seed location. We see that the performance is sensitive to the selection of the tracking plane. If the plane is not well-selected, the segmentation result can be empty as the example of tracking along the x-axis shows. The best result is tracking along the y-axis which is still worse than the proposed method since it considers potential turning points to leverage the 3D structure. The best result is tracking along the y-axis. However, it still falls short of the proposed method’s performance, as the latter takes into account potential turning points to exploit the 3D structure and perform long-term tracking.
TABLE IV: Ablation studies on different seed sampling strategies. The proposed recursive seed sampling approach achieves the best overall accuracy with comparable speed.
Strategy Pre. (%) Rec. (%) Acc. (%) Speed (sec)
@@
Dense sampling 86.94 24.84 23.95 2001($\uparrow$ 552%)
Recursive sampling 78.41 74.97 62.14 335 ($\uparrow$ 9%)
@@
Effect of the Recursive Redirection Strategy. To perform long-term tracking and fully leverage the 3D blood vessel structure, we introduced recursive seed sampling by considering potential turning points. To validate its effectiveness, we report the results on the BvEM-Macaque volume in Tab. IV where the runtime comparison for segmentation prediction on the entire BvEM-Macaque data is also included. Compared to the baseline without recursive seed sampling “naive” and dense seed sampling “dense”, the proposed method achieves the best performance with less than 10% running time increase. One naive baseline is to remove the recursive seed sampling component and not consider any potential turning points. This strategy is simple and fast but it fails to exploit the 3D shape prior leading to poor performance. Another strategy is to select the best plane for every tracking step/slice, which densely performs SAM segmentation on each step across three planes. Therefore, the performance is limited. The proposed method achieves the best performance with less than 10% running time increase on the BvEM-Macaque dataset. Unfortunately, it significantly increases the computation cost. Compared to our method, DenseSAM’s running time is 5.97 times longer since it needs to segment 3 planes for every tracking step. However, we were surprised to observe that the performance of the Dense Redirection strategy was even worse. This could be attributed to the frequent axis changes potentially leading to the omission of certain parts of the blood vessel and causing splitting errors.
@@
TABLE V: Ablation study results on the choice of SAM models. The MobileSAM model achieves better performance with faster inference speed and smaller model size.
Backbone Pre (%) Rec (%) Acc (%) Speed Parameters
SAM [15] 77.65 66.03 55.48 1535s 615M
MobileSAM [36] 78.41 74.97 62.14 335s 9.66M
@@
Quality Analysis. We delve deeper into the plane dynamics in Fig. 11 and 11. The majority of the selected plane predominantly tracks along the y-axis. This observation aligns with the experimental results presented in Tab.III, where it is evident that a single plane tracking with the y-axis outperforms the z-axis and x-axis. This is because the blood vessel flows mainly along the y-axis in the test volume. In Fig. 11, we explore the size dynamics using various methods, where the mean derivation is shown in parentheses. The size variation observed with our proposed method is relatively smaller compared to tracking along the y-axis and significantly less than when tracking along the z-axis and x-axis.
Other Ablations. To further improve the efficiency, we use MobileSAM instead of the original SAM in practice. As shown in Table V, the inference time of MobileSAM is 22% of the original SAM which confirms that MobileSAM significantly improves the inference speed. Moreover, the performance of MobileSAM is even better than the original SAM, possibly because the distilled small model is less prone to overfitting to the original natural image domain.Fig. 8 illustrates the performance with varying hyperparameters $\eta$ and $\tau$.
Figure 12: The failure cases of the proposed method. Three prediction and ground truth pairs are shown.
V-D Failure Cases